# Convolution estimates for the Boltzmann gain operator with hard-spheres

### v3

@@ 1.1. Problem setup and goal of this paper
where the advection operator $\partial_{t}+v\cdot\nabla_{x}$ accounts for the rectilinear motion of particles, while $Q(f,f)$ is the quadratic collisional operator encoding the binary interactions between gas particles. The collisional operator $Q(f,f)$ is given by
@@
where $v,v_{1}$ denote the pre-collisional velocities of the incoming particles, $v^{*}$, $v_{1}^{*}$ denote the post-collisional velocities, and $u:=v-v_{1}$ denotes the relative velocity of the particles before the collision. Simplifying the notation slightly, we drop the dependence of the arguments on $t,x$ since it is clearly implied by the context.
@@
  • The case $0<\gamma\leq 1$ corresponds to hard potentials, where the interaction is strong and of short range. They include the celebrated hard-sphere model ($\gamma=1,\,b=1$), where particles interact like impenetrable billiard balls. Hard-spheres are of particular importance since it was the model originally introduced by Boltzmann [8] and Maxwell [30] and remains the only model for which the Boltzmann has been rigorously derived from deterministic many particle dynamics [27, 22, 15].
@@
Both physically and mathematically, these three types of potentials lead to different kind of obstacles that typically require separate treatment. In this paper, we focus on hard potentials. In the course of the proof, we also improve existing results on Maxwell molecules.
Regarding the assumptions on the angular cross-section $b$, condition (H.2) corresponds to indistiguishability of the particles, while the boundedness condition (H.3) is a slightly stronger assumption than the usual Grad’s cut-off, which is the mere integrability condition $\int_{{\mathbb{S}}^{2}}b(\widehat{u}\cdot\sigma)\,d\sigma<\infty.$ Yet (H.3) covers most cases of physical interest for hard potentials, and in particular the classical hard-sphere model ($\gamma=1,\,b=1$). Most importantly, since (H.3) is a special case of Grad’s cut-off, it implies that the collisional operator can be written in gain-loss form
@@
Exploiting spherical symmetry, it is often convenient to restrict the collisional operators to the half-sphere. Namely, denoting $\chi:=\mathds{1}_{(0,+\infty)},$ one obtains
$Q(f,f)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)\Big{(}f(v^{*})f(v_{1}^{*})-f(v)f(v_{1})\Big{)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\hskip 28.45274pt+\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)\Big{(}f(v^{*})f(v_{1}^{*})-f(v)f(v_{1})\Big{)}\chi(-\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)\Big{(}f(v^{*})f(v_{1}^{*})-f(v)f(v_{1})\Big{)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\hskip 28.45274pt+\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(-\widehat{u}\cdot\sigma)\Big{(}f(v_{1}^{*})f(v^{*})-f(v)f(v_{1})\Big{)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)\Big{(}f(v^{*})f(v_{1}^{*})-f(v)f(v_{1})\Big{)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
where in the negative part we used the substitution $\sigma\to-\sigma$ (which just interchanges $v^{*}$ and $v_{1}^{*}$ by (1.3)) and for the last equality we used (H.2).
Motivated by this observation, we extend the collisional operators to bilinear forms in the following way:
@@
(1.12) $Q^{+}(f,g)$ $:=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})g(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
(1.13) $Q^{-}(f,g)$ $:=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v)g(v_{1})\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}.$
@@
In contrast, the gain operator $Q^{+}(f,g)$ has a much richer structure due to the averaging over the angular variable, and exhibits regularization effects [25, 28]. As a result, it is the main object to understand when developing the mathematical theory of the Boltzmann and related equations.
Quantifying these regularizing effects is a subtle task though due to the complex structure of the operator. The first step is to choose a suitable topology to measure the gain of regularity. The most relevant functional analytic setting is polynomially weighted Lebesgue spaces only in the velocity variable ${\langle}v{\rangle}^{-k}L^{p}_{v}$ for several reasons:
  • Such bounds are space invariant thus they can be applied whether the spatial domain is bounded or unbounded. Indeed, ${\langle}v{\rangle}^{-k}L^{q}_{x}L^{p}_{v}$ estimates easily follow by Hölder’s inequality. This feature is important to address questions related to the hydrodynamic limit of (1.1). In that field, most mathematical results are proved in spatially bounded domains such as the torus. Without being exhaustive, we refer to the recent work of Deng-Hani-Ma [16] for a complete derivation of fundamental fluid dynamics equations from many-body classical dynamics, which followed their groundbreaking work [15] on the long-time derivation of the Boltzmann equation from hard-sphere dynamics, that extended Lanford’s theorem [27, 22] globally in time.
  • Polynomial weights are the most suitable for studying the long-time behavior of the equation for large data, which is a major open problem, as outlined in [37]. Indeed, large data results for the space homogeneous (i.e. space independent) Boltzmann equation for hard potentials have been proved by establishing generation and propagation of moments [34, 20, 17, 38, 40, 11, 31], which correspond exactly to Lebesgue spaces with polynomial weights in $v$. This position was already advocated by L. Desvillettes and C. Villani [18], where the asymptotic behavior of solutions that remain bounded in such spaces is fully characterized. As noted by these authors, this setting is also outside of the scope of linearization techniques, which typically require exponential weights.
  • Working with polynomial weights is also more natural in the closely related field of wave turbulence, where equilibria are typically inverse polynomials in $v$ (e.g. Rayleigh-Jeans distributions or the celebrated Kolmogorov-Zakharov spectra, see e.g. Nazarenko [33] for more details). Therefore, ideas from this setting could be helpful in studying wave kinetic equations as well, as shown in our previous work [3]. In an upcoming work [5] we are planning to use ideas and techniques of the current paper in order to prove optimal local well-posedness of the wave kinetic equation in $L^{p}$ spaces for any $2\leq p\leq\infty$. Such a result would extend the work of Germain-Ionescu-Tran [23] who obtained optimal well-posedness in $L^{2}$ and in $L^{\infty}$.
@@
In this paper, we prove new moment-preserving polynomially weighted convolution estimates for $Q^{+}(f,g)$ with hard potentials, including hard-spheres. Our approach relies crucially on a novel cancellation mechanism dealing with pathological collisions that accumulate energy to only one of the outgoing particles, which we call energy-absorbing collisions. As we will see, this type of collisions distinguishes hard potentials from Maxwell molecules. Hence, a thorough understanding of these is crucial for the study of hard potentials.
@@
After the foundational works of Carleman [13, 14] ($1<p<\infty$) and Arkeryd [6] ($p=\infty$) on the propagation of $L^{p}$–norms of solutions to the homogeneous Boltzmann equation, it was Gustafsson [24, 25] who first noticed the convolution structure of the gain collisional operator by proving a Young-type inequality for $Q^{+}(f,g)$, using an intricate nonlinear interpolation procedure. Later, in a pioneering work, Lions [28] used Fourier integral operators techniques to show that $Q^{+}(f,g)$ gives a derivative gain. Wennberg [39] simplified the proof of Lions and arrived at the same estimates by using the Carleman representation and classical Fourier analysis tools. See also the works [12, 29] for another family of estimates on the quadratic operator $Q^{+}(f,f)$ that are obtained by much simpler means. Based on these results, Toscani-Villani [36] and Mouhot-Villani [32] studied the propagation of $L^{p}$–norms and smoothness of solutions to the homogeneous Boltzmann equation. However, all these early results were obtained assuming point-wise truncations in the angular variable, namely singular collisions such as grazing $(\widehat{u}\cdot\sigma=0)$ or frontal $(\widehat{u}\cdot\sigma=\pm 1)$ are discarded.
More recently, Alonso-Carneiro [1] used Fourier analysis and radial symmetrization techniques to obtain a Young-type inequality for Maxwell molecules and hard potentials, under certain integrability assumption on the cross-section $b$ that replaced point-wise truncations. The dependence on the cross-section is embedded in the constants of their estimates. Special cases of those estimates were previously obtained in [19] and [21]. Soon after that, Alonso-Carneiro-Gamba [2] extended the results of [1] to inelastic collisions, and also proved a Hardy-Littlewood-Sobolev type of inequality for soft potentials ($3<\gamma<0$).
@@
Simple moment increasing estimates of this kind can be obtained by treating the operator as a Maxwell molecule with increased weights in the arguments (see [1, 2]). However, these estimates unavoidably suffer the major drawback of being moment-increasing. Namely for $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$, the estimate provided by [1, 2] is of the form:
$\|{\langle}v{\rangle}^{k}\,Q^{+}_{cl}(f,g)\|_{L^{r}_{v}}\lesssim\|{\langle}v{\rangle}^{k+\gamma}\,f\|_{L^{p}_{v}}\|{\langle}v{\rangle}^{k+\gamma}\,g\|_{L^{q}_{v}},\quad{\langle}v{\rangle}:=\sqrt{1+|v|^{2}},$
which is not sufficient to prove local/global well-posedness of the Boltzmann equation in spaces of the form ${\langle}v{\rangle}^{-k}L^{q}_{x}L^{p}_{v}$, or propagation of moments. The main difficulty is to offset the linear growth of the potential. Following Lions [28], a now classical (but highly technical) approach relies on decay properties of oscillatory integrals. Indeed denoting $\widehat{f}$ and $\widehat{g}$ the Fourier transforms of $f$ and $g,$ (1.9) can be written (taking $b=1$ for simplicity)
@@
The decay in $|u|$ compensates for the growth of the potential. The remaining challenge is to prove $L^{p}$–boundedness of the resulting Fourier integral operator. A very recent paper [26] adapts this method to the hard potential case to produce a moment-preserving estimate for the gain operator with hard potentials. However, in the case of hard-spheres, this work only gives a single estimate, namely $\|{\langle}v{\rangle}^{k}Q^{+}_{cl}(f,g)\|_{L^{6}_{v}}\lesssim\|{\langle}v{\rangle}^{k}f\|_{L^{1}_{v}}\|{\langle}v{\rangle}^{k}g\|_{L^{2}_{v}}$.
In comparison, we propose a kinetic alternative to the Fourier analytic approach, which we now present. This approach allows us to prove a much wider class of estimates.
@@ 1.2. Heuristics
For simplicity, in this subsection we focus on the case of hard-spheres i.e. $\gamma=1$ and $b=1$, which is the prototype hard potential model. We write the weighted (1.9) as
${\langle}v{\rangle}^{k}Q^{+}(f,g)(v)=2\int_{\mathbb{R}^{3}\times{\mathbb{S}}^{2}}\frac{|u|{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{k}}f_{k}(v^{*})g_{k}(v_{1}^{*})\chi(\widehat{u}\cdot\sigma)\,d\sigma dv_{1}.$
@@
We distinguish two types of collisions: energy-distributing and energy-absorbing. Each of these requires different treatment.
@@
In most instances, energy is distributed among the outgoing particles i.e. both $|v^{*}|^{2}$ and $|v_{1}^{*}|^{2}$ are comparable to the energy $E$. We call these collisions energy-distributing. Since $|u|\leq|v|+|v_{1}|\lesssim E^{1/2}$, we can then easily estimate
@@
as long as $k\geq 1$. In other words, in the most likely scenario, the weights offset the growth of the potential and the operator behaves like a Maxwell molecule. For this reason, we provide an exhaustive treatment of Maxwell molecules, which is of independent interest. In fact, we improve and extend existing results for such operators. As mentioned, convolution estimates without additional angular cut-offs for Maxwell molecules have been proved in [1, 2]. In these works, though, the implicit constants in the estimates for $b=1$ are finite only when $r<3$. We opt for a self-contained, purely kinetic approach that allows us to extend the existing estimates on Maxwell molecules for $b$ satisfying (H.1)–(H.3) to arbitrary large $r$ (even $r=\infty$), as long as $p<3$. This is the content of Theorem 3.1 ($r=\infty$) and Theorem 3.2 ($r<\infty$). We stress that our technique is also simpler, and only relies on elementary kinetic tools. Moreover, most of the aforementioned papers that introduce angular cut-offs in the gain operator fall into this category, since the cut-offs allow only energy-distributing collisions to occur. Our method could potentially be used to revisit and extend these works as well.
@@
The second, more pathological, scenario consists of one of the outgoing particles accumulating almost all the energy of the collision. We call those energy-absorbing collisions. They distinguish hard potentials from Maxwell molecules. Following the billiard analogy, they can be visualized as two balls colliding, with one staying still after the collision while the other absorbs all of the energy.
@@
It is easy to see that energy-absorbing collisions happen with non-trivial probability. For instance, when $\sigma\approx\frac{v}{|v|}$, the collisional law (1.3) yields $v^{*}\approx 0,\quad v_{1}^{*}\approx v$. However, in this case $\mathcal{R}\approx|u|=|v^{*}-v_{1}^{*}|\approx|v|$, where for the second equality we used (1.6). This leads to a growth of moments which cannot be controlled point-wise. Hence a crude bound would result in a moment-increasing estimate, which is not sufficient to understand the time evolution of the system, even locally in time.

### v4

@@ 1.1. Problem setup and goal of this paper
where the advection operator $\partial_{t}+v\cdot\nabla_{x}$ accounts for the rectilinear motion of particles, while $Q(f,f)$ is the quadratic collisional operator encoding the binary interactions between gas particles. It is given by
@@
where $v,v_{1}$ denote the pre-collisional velocities of the incoming particles, $v^{*}$, $v_{1}^{*}$ denote the post-collisional velocities, and $u:=v-v_{1}$ denotes the relative velocity of the particles before the collision. Simplifying the notation slightly, we drop the dependence of the arguments on $t,x.$
@@
  • The case $0<\gamma\leq 1$ corresponds to hard potentials, where the interaction is strong and of short range. They include the celebrated hard-sphere model ($\gamma=1,\,b=1$), where particles interact like impenetrable billiard balls. Hard-spheres are of particular importance since it was the model originally introduced by Boltzmann [8] and Maxwell [30]. It remains the only model for which the Boltzmann has been rigorously derived from deterministic many particle dynamics [27, 22, 15].
@@
Both physically and mathematically, these three types of potentials lead to different kinds of difficulties that typically require separate treatment. In this paper we focus on hard potentials, i.e. we assume $0<\gamma\leq 1$. In the course of the proof, we also improve existing results on Maxwell molecules with bounded cross-sections.
Regarding the assumptions on the angular cross-section $b$, condition (H.2) corresponds to indistinguishably of the particles, while the boundedness condition (H.3) is a slightly stronger assumption than the usual Grad’s cut-off, which is the mere integrability condition $\int_{{\mathbb{S}}^{2}}b(\widehat{u}\cdot\sigma)\,d\sigma<\infty.$ Yet (H.3) covers most cases of physical interest for hard potentials, and in particular the classical hard-sphere model ($\gamma=1,\,b=1$). Most importantly, since (H.3) is a special case of Grad’s cut-off, it implies that the collisional operator can be written in gain-loss form
@@
Exploiting spherical symmetry, it is often convenient to restrict the collisional operators to a half-sphere. For hard potentials, the most natural restriction is to fix the outgoing particle with the largest velocity (for Maxwell molecules we will restrict to the half-sphere corresponding to $\widehat{u}\cdot\sigma>0$ instead, see Section 3). Namely, by conservation of energy we have $|v^{*}|^{2}+|v_{1}^{*}|^{2}=|v|^{2}+|v_{1}|^{2}:=E$, thus either $|v_{1}^{*}|^{2}\geq E/2$ or $|v^{*}|^{2}\geq E/2$. Since the set where $|v_{1}^{*}|^{2}=|v^{*}|^{2}=E/2$ is of measure zero, we obtain
$Q^{+}(f,f)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})f(v_{1}^{*})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1}$
$\hskip 28.45274pt+\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})f(v_{1}^{*})\mathds{1}_{|v^{*}|^{2}>E/2}\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})f(v_{1}^{*})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1}$
$\hskip 28.45274pt+\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(-\widehat{u}\cdot\sigma)f(v_{1}^{*})f(v^{*})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1}$
$=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})f(v_{1}^{*})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1},$
where in the second part of the summation we used the substitution $\sigma\to-\sigma$ (which just interchanges $v^{*}$ and $v_{1}^{*}$ by (1.3)) and for the last equality we used (H.2). Similarly, we obtain
(1.11) $Q^{-}(f,f)=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v)g(v_{1})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1}.$
Motivated by this computation, we introduce the bilinear forms:
@@
(1.13) $Q^{+}(f,g)$ $:=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v^{*})g(v_{1}^{*})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1},$
(1.14) $Q^{-}(f,g)$ $:=2\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}b(\widehat{u}\cdot\sigma)f(v)g(v_{1})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma\,dv_{1}.$
@@
In contrast, the gain operator $Q^{+}(f,g)$ has a much richer structure due to the averaging over the angular variable and exhibits regularization effects [25, 28]. As a result, it is the main object to understand when developing the mathematical theory of the Boltzmann and related equations.
Quantifying these regularizing effects is subtle, though, because of the complex structure of the operator. The first step is to choose a suitable topology to measure the gain of regularity. The most relevant functional analytic setting is polynomially weighted Lebesgue spaces only in the velocity variable ${\langle}v{\rangle}^{-k}L^{p}_{v},$ for several reasons:
  • Such bounds are space translation invariant, thus they can be applied whether the spatial domain is bounded or unbounded. Indeed, ${\langle}v{\rangle}^{-k}L^{q}_{x}L^{p}_{v}$ estimates easily follow from Hölder’s inequality. This feature is important to address questions related to the hydrodynamic limit of (1.1). In that field, most mathematical results are proved in spatially bounded domains such as the torus. Without being exhaustive, we refer to the recent work of Deng-Hani-Ma [16] for a complete derivation of fundamental fluid dynamics equations from many-body classical dynamics. This followed their groundbreaking work [15] on the long-time derivation of the Boltzmann equation from hard-sphere dynamics which extended Lanford’s theorem [27, 22] globally in time.
  • Polynomial weights are the most suitable for studying the long-time behavior of the equation for large data, which is a major open problem, as outlined in [37]. Large data results for the space homogeneous (i.e. space independent) Boltzmann equation for hard potentials have been proved by establishing generation and propagation of moments [34, 20, 17, 38, 40, 11, 31], which correspond exactly to Lebesgue spaces with polynomial weights in $v$. Moreover, as noted by L. Desvillettes and C. Villani [18], the evolution cannot be treated by linearization techniques which typically require exponential weights.
  • Working with polynomial weights is also more natural in the closely related field of wave turbulence, where equilibria are typically inverse polynomials in $v$ (e.g. Rayleigh-Jeans distributions or the celebrated Kolmogorov-Zakharov spectra, see e.g. Nazarenko [33] for more details). Therefore, ideas from the present paper could also be helpful in studying wave kinetic equations, as shown in our previous work [3]. In an upcoming article [5] we are planning to use ideas and techniques of the current paper in order to prove optimal local well-posedness of the wave kinetic equation in $L^{p}$ spaces for any $2\leq p\leq\infty$. Such a result would extend the work of Germain-Ionescu-Tran [23] who obtained optimal well-posedness in $L^{2}$ and in $L^{\infty}$.
@@
In this paper, we use the averaging present in the gain operator to prove new moment-preserving convolution estimates in such polynomially weighted spaces. We work in the case of hard potentials, and the results include hard-spheres. Our approach relies crucially on a novel cancellation mechanism to deal with pathological collisions that accumulate energy to only one of the outgoing particles. As we shall see, this type of collisions distinguishes hard potentials from Maxwell molecules. Hence, a thorough understanding of these interactions is crucial for the study of hard potentials.
@@
After the foundational works of Carleman [13, 14] ($1<p<\infty$) and Arkeryd [6] ($p=\infty$) on the propagation of $L^{p}$–norms of solutions to the homogeneous Boltzmann equation, it was Gustafsson [24, 25] who first noticed the convolution structure of the gain collisional operator by proving a Young-type inequality for $Q^{+}(f,g)$, using an intricate nonlinear interpolation procedure. Later, in a pioneering work, Lions [28] used Fourier integral operators techniques to show that $Q^{+}(f,g)$ gives a derivative gain. Wennberg [39] simplified the proof of Lions and arrived at the same estimates by using the Carleman representation and classical Fourier analysis tools. See also the works [12, 29] for another family of estimates on the quadratic operator $Q^{+}(f,f)$ that are obtained by much simpler means. Based on these results, Toscani-Villani [36] and Mouhot-Villani [32] studied the propagation of $L^{p}$–norms and smoothness of solutions to the homogeneous Boltzmann equation. However, all these early results were obtained assuming point-wise truncations in the angular variable. More precisely, in these works singular collisions such as grazing $(\widehat{u}\cdot\sigma=0)$ or frontal $(\widehat{u}\cdot\sigma=\pm 1)$ are discarded.
More recently, Alonso-Carneiro [1] used Fourier analysis and radial symmetrization techniques to obtain a Young-type inequality for Maxwell molecules, under integrability assumptions on the cross-section $b$ that replaced point-wise truncations. The dependence on the cross-section is embedded in the constants of their estimates. Special cases of those estimates were previously obtained in [19] and [21]. Soon after that, Alonso-Carneiro-Gamba [2] extended the results of [1] to inelastic collisions, and also proved a Hardy-Littlewood-Sobolev type of inequality for soft potentials ($3<\gamma<0$).
@@
Simple estimates of this kind can be straightforwardly obtained from the Maxwell molecule case, at the price of increasing weights on the arguments (see [1, 2]). These bounds are moment-increasing. More precisely for $\frac{1}{p}+\frac{1}{q}=1+\frac{1}{r}$, the estimate provided by [1, 2] is:
$\|{\langle}v{\rangle}^{k}\,Q^{+}(f,g)\|_{L^{r}_{v}}\lesssim\|{\langle}v{\rangle}^{k+\gamma}\,f\|_{L^{p}_{v}}\|{\langle}v{\rangle}^{k+\gamma}\,g\|_{L^{q}_{v}},\quad{\langle}v{\rangle}:=\sqrt{1+|v|^{2}},$
which is not sufficient to prove local/global well-posedness of the Boltzmann equation in spaces of the form ${\langle}v{\rangle}^{-k}L^{q}_{x}L^{p}_{v}$, or propagation of moments. The main difficulty is to offset the linear growth of the potential. Following Lions [28], a now classical (but highly technical) approach is based on decay properties of oscillatory integrals. Indeed, denoting $\widehat{f}$ and $\widehat{g}$ the Fourier transforms of $f$ and $g,$ (1.13) can be written (taking $b=1$ for simplicity)
@@
The decay in $|u|$ compensates for the growth of the potential. The remaining challenge is to prove $L^{p}$–boundedness of the resulting Fourier integral operator. A very recent paper [26] adapts this method to the hard potential case to produce a moment-preserving estimate for the gain operator with hard potentials. However, in the case of hard-spheres, this work only gives a single estimate, namely $\|{\langle}v{\rangle}^{k}Q^{+}(f,g)\|_{L^{6}_{v}}\lesssim\|{\langle}v{\rangle}^{k}f\|_{L^{1}_{v}}\|{\langle}v{\rangle}^{k}g\|_{L^{2}_{v}}$.
In comparison, we propose a kinetic alternative to the Fourier analytic approach, which we present now. It allows us to prove a much wider set of estimates.
@@ 1.2. Heuristics
For simplicity, in this subsection we focus on the case of hard-spheres i.e. $\gamma=1$ and $b=1$, which is the prototypical hard potential model. We write the weighted (1.13) as
${\langle}v{\rangle}^{k}Q^{+}(f,g)(v)=2\int_{\mathbb{R}^{3}\times{\mathbb{S}}^{2}}\frac{|u|{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{k}}f_{k}(v^{*})g_{k}(v_{1}^{*})\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\,d\sigma dv_{1}.$
@@
We distinguish two types of collisions: energy-distributing and energy-absorbing. Each of these requires a different argument.
@@
In most cases, energy is distributed among the outgoing particles, i.e. $|v^{*}|^{2}$ is comparable to the energy as well. We call these collisions energy-distributing. In Figure 1, we show an example of such a collision.
Figure 1. Energy-distributing collisions i.e. $|v^{*}|\approx|v_{1}^{*}|\approx E^{1/2}$
Now, since $|u|\leq|v|+|v_{1}|\lesssim E^{1/2}$, we can easily estimate
@@
as long as $k\geq 1$. In other words, in the most likely scenario, the weights offset the growth of the potential and the operator behaves like a Maxwell molecule. For this reason, we provide an exhaustive treatment of Maxwell molecules, which is of independent interest. In fact, we improve and extend existing results for such operators. As mentioned, convolution estimates without additional angular cut-offs for Maxwell molecules have been proved in [1, 2]. However, in these works the implicit constants in the estimates for $b=1$ are finite only when $r<3$. We opt for a self-contained, purely kinetic approach that allows us to extend the existing estimates on Maxwell molecules for $b$ satisfying (H.1)–(H.3) to arbitrarily large $r$ (even $r=\infty$), as long as $p<3$. This is the content of Theorem 3.1 ($r=\infty$) and Theorem 3.2 ($r<\infty$). We stress that our technique is also simpler, and only relies on elementary kinetic tools. Moreover, most of the aforementioned papers that introduce angular cut-offs in the gain operator fall into this category, since the cut-offs allow only energy-distributing collisions to occur. Our method could be used to revisit and extend these works as well.
@@
The second, more pathological scenario consists of one of the outgoing particles accumulating almost all the energy of the collision i.e. $|v_{1}^{*}|\approx E^{1/2}$, $v^{*}\approx 0$. We call those energy-absorbing collisions. They distinguish hard potentials from Maxwell molecules. Following the billiard analogy, they can be visualized as two balls colliding, one staying still after the collision, while the other absorbs all of the energy. In Figures 2, 3, we provide examples of the two prototype energy-absorbing collisions, depending on whether $|v_{1}|\gtrsim|v|$ or $|v_{1}|\ll|v|$.
Figure 2. Energy-absorbing collision with $|v_{1}|\gtrsim|v|$, $v^{*}_{1}\approx v+v_{1}$
Figure 3. Energy-absorbing collisions with $|v_{1}|\ll|v|$, $v_{1}^{*}\approx v$
@@
It is easy to see that energy-absorbing collisions happen with non-trivial probability. For example when $\sigma\approx\frac{2V}{|u|}$, the collisional law (1.3) yields $v^{*}\approx 0,\quad v_{1}^{*}\approx V$. However, in this case $\mathcal{R}\approx|u|=|v^{*}-v_{1}^{*}|\approx|v_{1}^{*}|\approx E^{1/2}$, where for the second equality we used (1.6). This leads to a growth of moments which cannot be controlled point-wise. Hence a crude bound would result in a moment-increasing estimate, which is not sufficient to understand the time evolution of the system, even locally in time.